OpenStudy (anonymous):

What is d/dx(10^x) ?

terenzreignz (terenzreignz):

Okay... you know what is \[\Large \frac{d}{dx}e^x=\color{red}?\]

OpenStudy (anonymous):

yep! - e^x itself.

terenzreignz (terenzreignz):

That is correct :) Now let's play with this expression a bit... \[\Large 10^x =[ e^{\ln 10}]^x\] Is this not so? :D

terenzreignz (terenzreignz):

Well, anyway, as per your laws of exponents... \[\Large 10^x = [e^{\ln 10}]^x=e^{(\ln10)x}\] So you may now use the chain rule... not forgetting that ln(10) is JUST A CONSTANT. \[\Large \frac{d}{dx}[e^{(\ln10)x}]=\color{red}?\]

OpenStudy (amistre64):

the general rule is:\[\frac d{dx}[a^x]=a^x~ln(a)\] when a=e, that simplifies to the e^x "rule"

terenzreignz (terenzreignz):

^ Well you're no fun... HAHA Just kidding Peace ^_^

terenzreignz (terenzreignz):

\[\Large \frac{d}{dx}[e^{(\ln10)x}]=e^{x\ln 10}\cdot \frac{d}{dx}(x\cdot \ln10)\] \[\Large = e^{\ln 10^x}\cdot \ln 10\] \[\Large = 10^x\cdot \ln 10\] Just as it should be :) Feel free to replace 10 with a XD

OpenStudy (amistre64):

:) i asked my teacher why there is a constant rule AND and a power rule: isnt kx^0 = k*0x^{-1} = 0? but since there is a "what if x=0" undefinablility .... they tweaked it

OpenStudy (anonymous):

Man, You're good, but to be honest, I didn't get much of what you said :'( - @terenzreignz

terenzreignz (terenzreignz):

Oh, I rushed it. Just for you to appreciate the awesomeness of simplicity :D I can go into more detail if you so desire :D

OpenStudy (anonymous):

And, @amistre64 , Thanks a lot for the help, but i want an explanation on how exactly does it happen!

terenzreignz (terenzreignz):

Here, let's go for the general case. \[\Large \frac{d}{dx}a^x\]

OpenStudy (anonymous):

Okay..

terenzreignz (terenzreignz):

we only know the more 'traditional' derivatives plus the fact that\[\Large \frac{d}{dx}e^x = e^x\] aye? :)

OpenStudy (anonymous):

Aye aye :P

terenzreignz (terenzreignz):

Okay. Then let's take a look at \(\large a^x\) For starters, \[\Large a = e^{\ln a}\] right?

terenzreignz (terenzreignz):

That part understood?

OpenStudy (anonymous):

Working on it ! :P

terenzreignz (terenzreignz):

I mean, it's part of the properties of logarithms... \[\Huge b^{\log_bx}=x\\\Huge e^{\ln \ x}=x\]

OpenStudy (anonymous):

Oohh kaay! So we're just supposed to mug up the property, no ? :P And, The rest is fine! Thank You! :)

terenzreignz (terenzreignz):

Are you sure you get it now? Well, the rest is there, just not in so much detail....